OpenStudy (anonymous):

For function: g(x) = (x – 8)^2/3 (a) Find the critical numbers. (Enter your answers from smallest to largest. Enter NONE in any unused answer blanks.) (b) Find the open intervals on which the function is increasing or decreasing. (If you need to use or –, enter INFINITY or –INFINITY, respectively. Enter NONE in any unused answer blanks.) Increasing:(or[ and )or] Decreasing:" and "

OpenStudy (anonymous):

find the equation of the normal to y=xsinx at x=pi/2

OpenStudy (anonymous):

can you show me process too.

OpenStudy (anonymous):

not enough for me

OpenStudy (anonymous):

how do you do

OpenStudy (anonymous):

oho!! sorry

OpenStudy (ℐℵk℣ⱱøƴȡ◆):

wat?

OpenStudy (solomonzelman):

Yeah, seriously... WHAT ??!!

OpenStudy (anonymous):

This came directly from my homework in calculus.

OpenStudy (solomonzelman):

By critical points the mean the restriction(s) for the domain ?

OpenStudy (solomonzelman):

In this case, since the denominator is odd, and all you get is, `∛ (x - 8) ²` you wouldn't have any domain restrictions, because with an oddth root, in this case cube root, you will never come up with an imaginary/complex number. You don't have any denominator that has a variable either.

OpenStudy (solomonzelman):

And you can see that when `x→±∞, y→±∞` (as opposed to `x→±∞, y→∓∞`) Right ?

OpenStudy (anonymous):

That looks the same x+-inf, y+-inf as opposed to x+- inf, y+- inf

OpenStudy (solomonzelman):

You don't have the correct graph. I am sorry to say this

OpenStudy (solomonzelman):

actually, hold on

OpenStudy (anonymous):

I put it in my graphing calc, it should be that one...

OpenStudy (solomonzelman):

well so far I can predict that it is either top left, or bottom right corner.

OpenStudy (anonymous):

but i also have no clue how to do the beginning portion of this problem.. all i know how to do is graph it in my calc..

OpenStudy (solomonzelman):

Do you know how to graph a line ones you are given y=mx+b ?

OpenStudy (anonymous):

no.. :(

OpenStudy (solomonzelman):

So, you can't graph (for example) y=4x+3 ?

OpenStudy (anonymous):

oh, i just put that in my calc in the y= function , yeah i can do that

OpenStudy (solomonzelman):

Noo..... can you graph `y=4x+3` by hand ?

OpenStudy (solomonzelman):

y=4x+3 let's plug in some random numbers for x, y=4(1)+3 y=4+3 y=7, So we know that when x=1, y=7. Thus we get the point `(1,7)` Can you derive couple points like I did just now? (By plugging in ANY number for x) ?

OpenStudy (anonymous):

y=4(2)+3 y=8+3 y=11 so then x=2 y=11

OpenStudy (anonymous):

\[g(x) = (x – 8)^{2/3}\] take the derivative and see where it is undefined that will be at \(x=8\)

OpenStudy (anonymous):

that is all,you can pretty much do it with your eyeballs